# From Gaze to Insight: Bridging Human Visual Attention and Vision Language Model Explanation for Weakly-Supervised Medical Image Segmentation

### v1

@@ Abstract
Medical image segmentation remains challenging due to the high cost of pixel-level annotations for training. In the context of weak supervision, clinician gaze data captures regions of diagnostic interest; however, its sparsity limits its use for segmentation. In contrast, vision-language models (VLMs) provide semantic context through textual descriptions but lack the explanation precision required. Recognizing that neither source alone suffices, we propose a teacher-student framework that integrates both gaze and language supervision, leveraging their complementary strengths. Our key insight is that gaze data indicates “where” clinicians focus during diagnosis, while VLMs explain “why” those regions are significant. To implement this, the teacher model first learns from gaze points enhanced by VLM-generated descriptions of lesion morphology, establishing a foundation for guiding the student model. The teacher then directs the student through three strategies: (1) Multi-scale feature alignment to fuse visual cues with textual semantics; (2) Confidence-weighted consistency constraints to focus on reliable predictions; (3) Adaptive masking to limit error propagation in uncertain areas. Experiments on the Kvasir-SEG, NCI-ISBI, and ISIC datasets show that our method achieves Dice scores of 80.78%, 80.53%, and 84.22%, respectively—improving 3–5% over gaze baselines without increasing the annotation burden. By preserving correlations among predictions, gaze data, and lesion descriptions, our framework also maintains clinical interpretability. This work illustrates how integrating human visual attention with AI-generated semantic context can effectively overcome the limitations of individual weak supervision signals, thereby advancing the development of deployable, annotation-efficient medical AI systems. Code is available at: https://github.com/jingkunchen/FGI.git.
@@
Weakly Supervised Segmentation, Human Visual Attention, Vision Language Model, Medical Image Analysis.
@@ 1 Introduction
Efforts to reduce annotation overhead have sparked interest in weakly supervised segmentation, where coarse labels such as image-level tags, scribbles, bounding boxes, or points guide model training [4, 5, 6, 7]. Although these forms of supervision offer a more efficient alternative to full manual delineation, they frequently lack the precise spatial details necessary to capture subtle pathological boundaries or complex anatomical structures [8]. Consequently, the community has continued to explore additional weak signals or novel labeling strategies that strike a better balance between annotation economy and segmentation quality.
@@
  • We present the first weakly supervised framework that jointly leverages expert gaze patterns and vision-language model outputs for medical image segmentation, effectively combining human attention cues with rich semantic descriptions to overcome limitations of individual weak supervision signals.
@@ 3 Methodology
In this work, we introduce a teacher–student framework that harnesses the complementary strengths of human gaze and vision–language model (VLM) outputs for segmentation in medical imaging. As detailed in Section 2, gaze data can reliably indicate diagnostically relevant regions but often lack the spatial precision needed for accurate delineation. In contrast, VLMs provide high-level semantic context but may not localize lesions precisely when used directly on medical images. To reconcile these aspects, our method integrates three complementary inputs: (i) Sparse Pseudo-Masks—capturing high-confidence gaze fixations; (ii) Comprehensive Pseudo-Masks—providing broader yet noisier lesion coverage; (iii) VLM-Derived Textual Embeddings—encoding morphological and semantic attributes.
As illustrated in Figure 1, our method begins with the teacher stage, where image features are enhanced with textual cues derived from a vision–language model. A fusion module then uses sparse pseudo-masks to supervise feature refinement via a partial cross‑entropy loss, yielding robust cross‑modal representations. The teacher model extracts these representations to guide the student network, which is trained on comprehensive pseudo‑masks via feature‑level knowledge distillation combined with a confidence‑aware consistency regularization strategy. Next, we provide a detailed description of each module within our framework.
Figure 1: Overall Pipeline of Our Teacher–Student Approach. (a) A large vision–language model is prompted with a structured request for shape, boundary details, and approximate coverage to generate textual descriptions. (b) The Teacher fuses these textual cues with sparse gaze pseudo-masks via partial cross-entropy loss to extract robust cross-modal features. (c) The Student learns from comprehensive pseudo-masks through knowledge distillation and confidence-aware consistency, while a Disagreement-Aware Random Masking strategy is applied to mitigate the impact of inconsistent annotations.
@@ 3.1 Obtaining Gaze Pseudo-Masks and Vision–Language Text
Sparse vs. Comprehensive Pseudo-Masks
@@
    Sparse Pseudo-Mask ($M_{\mathrm{sp}}$): a small, meticulously confirmed masks set capturing core diagnostic fixations, yielding precise but partial annotations.
@@
    Comprehensive Pseudo-Mask ($M_{\mathrm{comp}}$): a more extensive record covering potential lesion regions, yet containing noise (e.g., ambiguous or off-target gaze).
The sparse mask offers near-certain ground truth for a subset of pixels, while the comprehensive mask—though broader—risks overlabeling. Our method first capitalizes on the high certainty of $M_{\mathrm{sp}}$ in a teacher stage, then transitions to $M_{\mathrm{comp}}$ during the student stage.
@@
Since the sparse pseudo-mask $M_{\mathrm{sp}}$ provides supervision only on a subset of pixels, we employ a partial cross-entropy loss defined as:
$\mathcal{L}_{pCE}(y,\hat{y})$ $=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\,\log\!\bigl{(}\hat{y}_{i,c}\bigr{)},$ (8)
$\text{where}\quad y_{i,c}$ =$=\begin{cases}1,&\text{if pixel }i\text{ is labeled as class }c\text{ in }M_{\mathrm{sp}},\\[3.0pt] 0,&\text{otherwise.}\end{cases}$
@@ 3.2 Student Network: Feature-Level Knowledge Distillation and Confidence-aware Weighted Consistency
While the teacher focuses on partial but precise information, the student must handle a comprehensive pseudo-mask $M_{\mathrm{comp}}$, encompassing full lesion coverage (often with noise). We ensure the student maintains the teacher’s cross-modal insights through two key strategies.
@@
[…]
@@
Comprehensive Cross-Entropy & Random Masking.
Finally, we include a standard cross-entropy term over $M_{\mathrm{comp}}$. Splitting it across two lines to avoid overflow:
$\mathcal{L}_{\mathrm{CE}}$ $(M_{\mathrm{comp}})=-\sum_{p}\Bigl{[}M_{\mathrm{comp}}(p)\,\log\!\bigl{(}\widehat{Y}_{\mathrm{student}}(p)\bigr{)}$ (13)
$~{}+~{}\bigl{(}1-M_{\mathrm{comp}}(p)\bigr{)}\,\log\!\bigl{(}1-\widehat{Y}_{\mathrm{student}}(p)\bigr{)}\Bigr{]}.$
Combining it all,
$\mathcal{L}_{\mathrm{student}}=\mathcal{L}_{\mathrm{CE}}(M_{\mathrm{comp}})+\lambda_{\mathrm{AFC}}\,\mathcal{L}_{\mathrm{AFC}}+\lambda_{\mathrm{CWC}}\,\mathcal{L}_{\mathrm{CWC}}.$ (14)
Since some $M_{\mathrm{comp}}$ fixations might be blatantly incorrect, we also randomly zero out local patches where teacher–student masks strongly disagree—employing a Disagreement-Aware Random Masking (DARM) strategy—which selectively occludes regions with the inconsistency to force the student model to rely on more reliable contextual cues, reduce the impact of inconsistenct annotations, and simultaneously encourage exploration of less certain regions.
@@ 3.3 Rationale and Benefits
  • Sparse Pseudo-Masks: Reliable yet partial gaze annotations that can be used with partial cross-entropy, preventing over-penalization in unlabeled areas.
@@
  • Comprehensive Pseudo-Masks: Wider lesion coverage, albeit noisier, refined through teacher distillation and high-confidence consistency checks.
@@ 4.3 Performance on Kvasir-SEG (Endoscopic Polyp Segmentation)
Table 1: Quantitative comparison on the Kvasir-SEG dataset. Underline indicates the best result among weakly supervised methods.
@@ 4.4 Performance on NCI-ISBI (MRI Prostate Segmentation)
Table 2: Quantitative comparison on the NCI-ISBI dataset. Bold indicates the best result among weakly supervised methods. AT denotes annotation time.
@@ 4.5 Performance on ISIC (Skin Lesion Segmentation)
Table 3: Quantitative comparison on the ISIC dataset. Bold indicates the best result among gaze methods.
@@
UNet Full $80.97\pm 0.05$ $88.17\pm 0.10$ $7.96\pm 0.34$ $3.22\pm 0.10$
nnUNet Full $81.98\pm 0.29$ $88.96\pm 0.18$ $7.70\pm 0.20$ $2.98\pm 0.09$
@@ 4.6.4 Effect of Disagreement-Aware Random Masking
Table 7: Impact of the Disagreement-Aware Random Masking on segmentation performance (Kvasir-SEG). Dice scores are reported in %.
@@
       Without DARM        $\text{76.21}\pm\text{0.40}$
       With DARM        $\text{80.78}\pm\text{0.11}$
@@ 4.6.5 Effect of Cross-Modal Aggregation Depth
Table 8: Effect of the number of fusion stages in the cross-modal aggregation module on segmentation performance (Kvasir-SEG). Dice scores are reported in %.
@@
Figure 3: Qualitative visualization of segmentation outputs from gaze-based annotations. Red overlays highlight gaze points and boundaries for direct method comparison. LVMs’ image interpretation is displayed in the rightmost panel.
@@ 6 Conclusion
In this paper, we introduced a novel teacher–student framework that leverages two complementary weak supervisory signals—human gaze and vision–language model outputs—to achieve robust medical image segmentation under limited annotation conditions. Our framework first refines sparse yet highly reliable gaze pseudo-masks by fusing them with rich, text-derived morphological and contextual cues in a teacher network, which in turn distills these refined cross-modal features to guide a student network. The student model, trained on more comprehensive but noisier gaze data, is further optimized with confidence-aware consistency loss and a Disagreement-Aware Random Masking strategy that selectively occludes regions with disagreement region, encouraging exploration of less certain areas while reducing reliance on spurious annotations. Extensive experiments on three benchmark datasets (Kvasir-SEG, NCI-ISBI, and ISIC) demonstrate that our approach significantly narrows the performance gap to fully supervised methods while drastically reducing annotation costs. Moreover, the integration of semantic embeddings not only enhances segmentation accuracy but also improves interpretability, which is critical in clinical applications. Future work will extend this framework to additional imaging modalities and incorporate even richer textual or meta-clinical data, further streamlining annotation processes and bolstering model explainability in real-world medical settings.

### v2

@@ Abstract
Medical image segmentation remains challenging due to the high cost of pixel-level annotations for training. In the context of weak supervision, clinician gaze data captures regions of diagnostic interest; however, its sparsity limits its use for segmentation. In contrast, vision-language models (VLMs) provide semantic context through textual descriptions but lack the explanation precision required. Recognizing that neither source alone suffices, we propose a teacher-student framework that integrates both gaze and language supervision, leveraging their complementary strengths. Our key insight is that gaze data indicates “where” clinicians focus during diagnosis, while VLMs explain “why” those regions are significant. To implement this, the teacher model first learns from gaze points enhanced by VLM-generated descriptions of lesion morphology, establishing a foundation for guiding the student model. The teacher then directs the student through three strategies: (1) Multi-scale feature alignment to fuse visual cues with textual semantics; (2) Confidence-weighted consistency constraints to focus on reliable predictions; (3) Adaptive masking to limit error propagation in uncertain areas. Experiments on the Kvasir-SEG, NCI-ISBI, and ISIC datasets show that our method achieves Dice scores of 80.78%, 80.53%, and 84.22%, respectively—improving 3–5% over gaze baselines without increasing the annotation burden. By preserving correlations among predictions, gaze data, and lesion descriptions, our framework also maintains clinical interpretability. This work illustrates how integrating human visual attention with AI-generated semantic context can effectively overcome the limitations of individual weak supervision signals, thereby advancing the development of deployable, annotation-efficient medical AI systems. Code is available at: https://github.com/jingkunchen/FGI.
@@
Weakly Supervised Segmentation, Visual Attention, Vision Language Model, Medical Image Analysis
@@ 1 Introduction
Efforts to reduce annotation overhead have sparked interest in weakly supervised segmentation, with model training steered by coarse labels such as image-level tags, scribbles, bounding boxes, or points [4, 5, 6, 7, 8, 9]. Although these forms of supervision offer a more efficient alternative to full manual delineation, they frequently lack the precise spatial details necessary to capture subtle pathological boundaries or complex anatomical structures [10]. Consequently, the community has continued to explore additional weak signals or novel labeling strategies that strike a better balance between annotation economy and segmentation quality.
@@
  • We introduce the first weakly supervised framework for medical image segmentation. It fuses expert gaze patterns with vision–language model outputs to generate clinically precise segmentation cues. These cues are directly optimized via a partial cross‑entropy loss in our teacher network. By doing so, we overcome the limitations of individual weak supervision signals and elevate the VLM from a passive feature extractor to an active supervisory oracle.
@@ 3 Methodology
In this work, we introduce a teacher–student framework that harnesses the complementary strengths of human gaze and vision–language model (VLM) outputs for segmentation in medical imaging. As detailed in Section 2, gaze data can reliably indicate diagnostically relevant regions but often lack the spatial precision needed for accurate delineation. In contrast, VLMs provide high-level semantic context but may not localize lesions precisely when used directly on medical images. To reconcile these aspects, our method integrates three complementary inputs: (i) High-confidence Mask—capturing high-confidence gaze fixations; (ii) Broad-Coverage Masks—providing broader yet noisier lesion coverage; (iii) VLM-Derived Textual Embeddings—encoding morphological and semantic attributes.
As illustrated in Figure 1, our method begins with the teacher stage, where image features are enhanced with textual cues derived from a vision-language model. A fusion module then uses high-confidence mask to supervise feature refinement via a partial cross-entropy loss, yielding robust cross-modal representations. The teacher model extracts these representations to guide the student network, which is trained on broad-coverage masks via feature-level knowledge distillation combined with a confidence-aware consistency regularization strategy. Figure 2 summarizes all optimization terms used in our framework: the teacher is supervised by partial cross-entropy ($\mathcal{L}_{\mathrm{pCE}}$) on $\mathbf{M}_{\mathrm{hc}}$; the student is optimized with cross-entropy ($\mathcal{L}_{\mathrm{CE}}$) on $\mathbf{M}_{\mathrm{bc}}$ under disagreement-aware random masking (DRAM); cross-model transfer is enforced via Angular Feature Consistency ($\mathcal{L}_{\mathrm{AFC}}$), and reliability is promoted by Confidence-aware Weighted Consistency ($\mathcal{L}_{\mathrm{CWC}}$). Next, we provide a detailed description of each module within our framework.
[…]
Figure 2: Overview of the losses used in our framework. Left: Angular Feature Consistency ($\mathcal{L}_{\mathrm{AFC}}$) aligns intermediate features between the teacher and the student. Middle: Confidence-aware Weighted Consistency ($\mathcal{L}_{\mathrm{CWC}}$) regularizes predictions on confident positive/negative regions. Right: Supervision terms—the teacher uses partial cross-entropy ($\mathcal{L}_{\mathrm{pCE}}$) on high-confidence mask ($\mathbf{M}_{\mathrm{hc}}$), while the student uses cross-entropy ($\mathcal{L}_{\mathrm{CE}}$) on broad-coverage masks ($\mathbf{M}_{\mathrm{bc}}$) with disagreement-aware random masking (DRAM).
@@ 3.1 Obtaining Gaze Pseudo-Masks and Vision–Language Text
High-Confidence vs. Broad-Coverage Pseudo-Masks
@@
    High-Confidence Mask ($M_{\mathrm{hc}}$): a small, meticulously confirmed masks set capturing core diagnostic fixations, yielding precise but partial annotations.
@@
    Broad-Coverage Mask ($M_{\mathrm{bc}}$): a more extensive record covering potential lesion regions, yet containing noise (e.g., ambiguous or off-target gaze).
The high-confidence mask offers near-certain ground truth for a subset of pixels, while the broad-coverage mask—though broader—risks overlabeling. Our method first capitalizes on the high certainty of $M_{\mathrm{hc}}$ in a teacher stage, then transitions to $M_{\mathrm{bc}}$ during the student stage.
@@
Since the high-confidence mask $M_{\mathrm{hc}}$ provides supervision only on a subset of pixels, we employ a partial cross-entropy loss defined as:
[…]
@@ 3.2 Student Network: Feature-Level Knowledge Distillation and Confidence-aware Weighted Consistency
While the teacher focuses on partial but precise information, the student must handle a broad-coverage mask $M_{\mathrm{bc}}$, encompassing full lesion coverage (often with noise). We ensure the student maintains the teacher’s cross-modal insights through two key strategies.
@@
  • $\langle\cdot,\cdot\rangle$ denotes the channel-wise inner product,
  • $\|\cdot\|$ represents the $L^{2}$-norm computed along the channel dimension,
@@
Comprehensive Cross-Entropy & Random Masking
Finally, we include a standard cross-entropy term over $M_{\mathrm{bc}}$. Splitting it across two lines to avoid overflow:
$\mathcal{L}_{\mathrm{CE}}$ $(M_{\mathrm{bc}})=-\sum_{p}\Bigl[M_{\mathrm{bc}}(p)\,\log\!\bigl(\widehat{Y}_{\mathrm{student}}(p)\bigr)$ (13)
$~+~\bigl(1-M_{\mathrm{bc}}(p)\bigr)\,\log\!\bigl(1-\widehat{Y}_{\mathrm{student}}(p)\bigr)\Bigr].$
Since some pixels in the broad-coverage mask $M_{bc}$ can be unreliable, we apply Disagreement-Aware Random Masking (DARM) before computing $\mathcal{L}_{CE}(M_{bc})$: let $\widehat{\mathbf{Y}}_{\mathrm{T}},\widehat{\mathbf{Y}}_{\mathrm{S}}\in[0,1]^{C}$ denote the Teacher/Student supervision, and define the high-disagreement set $\mathcal{D}=\{\,p\mid\lVert\widehat{\mathbf{Y}}_{\mathrm{T}}(p)-\widehat{\mathbf{Y}}_{\mathrm{S}}(p)\rVert_{\infty}\geq\tau_{\mathrm{dis}}\,\}$ with $\tau_{\mathrm{dis}}\in(0,1)$. At each iteration, we sample non-overlapping $s\times s$ patches $\mathcal{P}\subseteq\mathcal{D}$ at rate $\rho$ and form a masked supervision $\tilde{M}_{bc}=M_{bc}\odot(1-\mathbf{1}_{\mathcal{P}})$, where $\odot$ denotes element-wise multiplication and $\mathbf{1}_{\mathcal{P}}(p)=1$ if $p$ lies in any selected patch (else 0). The cross-entropy is then computed on $\tilde{M}_{bc}$, i.e., $\mathcal{L}_{CE}(\tilde{M}_{bc})$, which selectively occludes inconsistent supervision, reduces the impact of noisy annotations, and encourages reliance on more reliable contextual cues. Combining the terms, the Student objective is.
$\mathcal{L}_{\mathrm{student}}=\mathcal{L}_{CE}(M_{bc})+\lambda_{\mathrm{AFC}}\,\mathcal{L}_{AFC}+\lambda_{\mathrm{CWC}}\,\mathcal{L}_{\mathrm{CWC}}.$ (14)
Here, $\mathcal{L}_{CE}(M_{bc})$ supervises the mask branch using the broad-coverage pseudo-label (implemented on its DARM-masked variant $\tilde{M}_{bc}$); $\mathcal{L}_{AFC}$ aligns Teacher–Student intermediate features at fusion stages $k{=}1\ldots 4$; and $\mathcal{L}_{\mathrm{CWC}}$ imposes confidence-aware consistency on the class-wise logits over $(\Omega_{\mathrm{pos}},\Omega_{\mathrm{neg}})$. The weights $\lambda_{\mathrm{AFC}}$ and $\lambda_{\mathrm{CWC}}$ balance the respective contributions. For completeness, the Teacher is optimized with the Partially Cross-Entropy loss $\mathcal{L}_{pCE}(M_{hc})$.
@@ 3.3 Rationale and Benefits
  • High-Confidence Mask: Reliable yet partial gaze annotations that can be used with partial cross-entropy, preventing over-penalization in unlabeled areas.
@@
  • Broad-Coverage Masks: Wider lesion coverage, albeit noisier, refined through teacher distillation and high-confidence consistency checks.
@@ 4.3 Performance on Kvasir-SEG (Endoscopic Polyp Segmentation)
Table 1: Quantitative comparison on the Kvasir-SEG dataset. Underline indicates the best result among weakly supervised methods. AT denotes annotation time.
@@ 4.4 Performance on NCI-ISBI (MRI Prostate Segmentation)
Table 2: Quantitative comparison on the NCI-ISBI dataset. Underline indicates the best result among weakly supervised methods. AT denotes annotation time.
@@ 4.5 Performance on ISIC (Skin Lesion Segmentation)
Table 3: Quantitative comparison on the ISIC dataset. Underline indicates the best result among gaze methods.
@@
UNet Full $0.97\pm 0.05 88$$8.17\pm 0.10$ $7.96\pm 0.34$ $3.22\pm 0.10$
nnUNet Full $1.98\pm 0.29 88$$8.96\pm 0.18$ $7.70\pm 0.20$ $2.98\pm 0.09$
@@ 4.6.3 Effect of Loss Components in the Student Model
Figure 4: Qualitative comparison on three benchmarks. Rows correspond to Kvasir-SEG, NCI-ISBI, and ISIC. Columns (left to right): Ground Truth (GT), Gaze Map (GM), UNet, TransUNet (TrUNet), nnUNet, GazeMedSeg (GMS), Ours, and the VLM-derived Description. Red overlays indicate gaze fixations and reference boundaries; predicted masks are rendered as semi-transparent overlays for visual comparison. Across representative cases, our method yields more complete lesion coverage and sharper boundary adherence while reducing spurious activations in distractor regions.
@@ 4.6.4 Effect of Disagreement-Aware Random Masking
Table 7: Impact of the disagreement-aware random masking on segmentation performance. Dice scores are reported in %.
@@
   w/o DARM    $\text{76.21}\pm\text{0.40}$
   w DARM    $\text{80.78}\pm\text{0.11}$
@@ 4.6.5 Effect of Cross-Modal Aggregation Depth
Table 8: Effect of the number of fusion stages in the cross-modal aggregation module on segmentation performance. Dice scores are reported in %.
@@ 6 Conclusion
In this paper, we introduced a novel teacher–student framework that leverages two complementary weak supervisory signals—human gaze and vision–language model outputs—to achieve robust medical image segmentation under limited annotation conditions. Our framework first refines sparse yet highly reliable gaze pseudo-masks by fusing them with rich, text-derived morphological and contextual cues in a teacher network, which in turn distills these refined cross-modal features to guide a student network. The student model, trained on more comprehensive but noisier gaze data, is further optimized with confidence-aware consistency loss and a Disagreement-Aware Random Masking strategy that selectively occludes regions with disagreement region, encouraging exploration of less certain areas while reducing reliance on spurious annotations. Extensive experiments on three benchmark datasets (Kvasir-SEG, NCI-ISBI, and ISIC) demonstrate that our approach significantly narrows the performance gap to fully supervised methods while drastically reducing annotation costs. Moreover, the integration of semantic embeddings not only enhances segmentation accuracy but also improves interpretability, which is critical in clinical applications. Despite these strengths, our method has limitations: it relies on the quality and coverage of clinician gaze data and the fidelity of vision–language model outputs—both of which can suffer from noise, domain gaps, or rare pathology appearances—and introduces additional computational complexity in the cross-modal fusion and distillation stages. Clinical deployment also requires further prospective validation and calibration of uncertainty estimates. Future work will extend this framework to additional imaging modalities, other forms of weak supervision (scribbles or automated boxes) and incorporate even richer textual or meta-clinical data, further streamlining annotation processes and bolstering model explainability in real-world medical settings. We also plan to explore active and human-in-the-loop strategies to selectively acquire gaze annotations, methods to better quantify and propagate uncertainty, lightweight fusion architectures to reduce compute overhead, and improved calibration of the language-derived semantic cues.
Acknowledgment
The authors gratefully acknowledge Dr. Tao Tan (Faculty of Applied Sciences, Macao Polytechnic University, Macao SAR, China) for thoughtful critiques and helpful discussions that strengthened the presentation and analysis of this paper. This work was supported in part by National Natural Science Foundation of China (62441235) and Beijing Natural Science Foundation (L257005).